# Harnessing Consistency for Robust Test-Time LLM Ensemble

### v1

@@ Abstract
Zhichen Zeng1  Qi Yu1  Xiao Lin1  Ruizhong Qiu1  Xuying Ning1 Tianxin Wei1Yuchen Yan1Jingrui He1Hanghang Tong1 1University of Illinois Urbana-Champaign {zhichenz, htong}@illinois.edu
@@ 1 Introduction
Extensive efforts on test-time LLM ensemble can be broadly categorized into two categories: token-level and response-level ensemble. Token-level ensemble aligns and fuses the token probabilities of different LLMs at each decoding step Yu et al. (2024); Yao et al. (2024); Huang et al. (2024); Xu et al. (2024c), enabling fine-grained real-time correction for each token generation. Response-level ensemble offers a more coarse-level ensemble by selecting either a complete response Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024) or a span Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024) from candidate outputs. Despite their success in improving ensemble quality, existing approaches largely overlook ensemble robustness against noisy or erroneous signals. For instance, incorrect token alignments can result in faulty probability fusion, while errors in model predictions may further compromise the correctness of ensemble outputs. Therefore, it is crucial for ensemble methods to detect and mitigate such potential errors during inference to ensure reliable performance.
@@
To bridge this gap, we propose a plug-and-play technique named CoRE to enhance both robustness and performance of LLM ensemble. Our preliminary observations reveal that ensemble failures are closely tied to large discrepancies between the token distributions of different LLMs. Motivated by this, we propose to harness both token and model consistency to achieve robust ensemble. At the token level, significant disparities in output probabilities for a specific token indicate fine-grained uncertainty, often stemming from token space misalignment. To address this, we introduce token consistency, which measures the disparity between each model’s token probability and a reference probability, serving as a low-pass filter to amplify reliable tokens and suppress inconsistent ones. At the model level, large divergence among full token distributions reflects model conflicts. To capture this, we define model consistency that promotes model outputs with high self-confidence and low divergence from peers, thereby strengthening consistent models and downweighting unreliable ones. A key advantage of the proposed CoRE is that it is orthogonal to various token-level ensemble methods, and thus can be seamlessly integrated with no additional inference cost.
We summarize the main contributions as follows
  • We are the first to investigate the robustness issue in LLM ensemble.
  • We assess consistency at both the token and model levels to enhance ensemble performance and robustness. This method can be seamlessly integrated with various LLM ensemble strategies to enable test-time ensemble correction with no additional cost.
  • We conduct extensive experiments across diverse benchmark tasks, model combinations, and ensemble methods. As shown in Figure 2, CoRE consistently improves baseline ensemble methods, achieving an average performance gain of 1.3% and 2.8% on Top-2 and Top-3 model ensemble, respectively.
Figure 2: Ensemble performance across six benchmarks. We report the average performance of different ensemble methods with (CoRE) and without (Vanilla) CoRE.
@@ 2.1 Test-time LLM Ensemble
Response-level Ensemble selects the most promising response from the multiple outputs to ensemble LLMs at a coarser-grained level. One line of work selects or synthesizes a full response among model outputs by either training-free approaches, such as perplexity scoring or majority voting Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024), or training-based approaches that learn to rank responses Si et al. (2023). Another line of work Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024) proposes ensembling at the span level, aiming to strike a balance between fine-grained correction and context-aware decision-making.
@@ 2.2 Model Consistency
Multi-model Consistency combines outputs from different LLMs through majority voting Trad and Chehab (2024); Niimi (2025) or collaborative reasoning Wang et al. (2024); Liang et al. (2023); Li et al. (2024b). Despite improving robustness, most focus on response-level agreement, overlooking fine-grained token-level consistency essential for detecting subtle errors in real time.
@@ 3 Methodology
In this section, we present our proposed CoRE for trustworthy LLM ensemble. We begin by introducing the notations and formally defining the LLM ensemble problem in Section 3.1. Then, in Section 3.2, we reveal a strong correlation between ensemble performance and consistency scores at both the token and model levels. Motivated by these findings, we introduce our token-level consistency measure in Section 3.3 and our distribution-level consistency measure in Section 3.4, both designed to improve ensemble performance and reliability.
@@ 3.1 Problem Formulation
Formally, we are given a main model with vocabulary set $\mathcal{V}_{\text{main}}$ and $N$ assistant models with vocabulary set $\mathcal{V}_{\text{assist}_{i}}$ with $i\in\{1,2,...,N\}$. We denote the predicted token distributions of main and assistant models as $\bm{p}_{\text{main}}$ and $\bm{p}_{\text{assist}_{i}}$, respectively. LLM ensemble first learns a token alignment matrix $\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{assist}_{i}}|\times|\mathcal{V}_{\text{main}}|}$ between the main model and the $i$-th assistant model, where $\bm{A}_{i}(v,u)=1$ indicates that token $v\in\mathcal{V}_{\text{main}}$ is aligned with token $u\in\mathcal{V}_{\text{assist}_{i}}$. Token prediction probabilities can be further projected into the main model token space via $\bm{p}_{\text{assist}_{i}}\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{main}}|}$. For simplicity, we denote the aligned probability as $\tilde{\bm{p}}_{\text{assist}_{i}}=\bm{p}_{\text{assist}_{i}}\bm{A}_{i}$. The aligned probabilities can be further ensembled via
@@ 3.2 Observations
First, we examine how token disparity, measure by the difference between the aligned probability and the reference probability on a specific token $v$, i.e., $\bm{\delta}_{i}(v)=|\tilde{\bm{p}}_{\text{assist}_{i}}(v)-\bm{p}^{*}(v)|$, reflects the alignment correctness. As shown in Figure 4(a), aligned tokens concentrate at lower disparities, while misaligned tokens exhibit a heavier right tail. A one-sided test of $H_{0}:\mathbb{E}_{v\in\text{misalign}}\bm{\delta}_{i}(v)\leq\mathbb{E}_{v\in\text{align}}\bm{\delta}_{i}(v)$ yields a $p$-value of $2.7\times 10^{-44}$, indicating that aligned tokens indeed have smaller disparity. We summarize this in Observation 1.
@@
Secondly, we evaluate how model confidence, measured by the entropy of the token probability $H(\tilde{\bm{p}}_{\text{assist}_{i}})$, relates to the answer correctness. As shown in Figure 4(b), correct answers exhibit lower entropy than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[H(\tilde{\bm{p}})|\text{wrong}]\leq\mathbb{E}[H(\tilde{\bm{p}})|\text{correct}]$ yields a $p$-value of $7.7\times 10^{-108}$, confirming that wrong answers have larger entropy than correct ones. We summarize this into Observation 2.
@@
Thirdly, we evaluate how model consistency, measured by the sum of RBF-transformed token disparities, i.e., $C_{\text{assist}_{i}}(v)=\allowbreak\sum_{v\in\mathcal{V}_{\text{main}}}\exp(-\bm{\delta}_{\text{assist}_{i}}(v)/\sigma)$, varies between correct and wrong answers. As shown in Figure 4(c), correct answers exhibit higher model consistency than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[C_{\text{assist}_{i}}|\text{wrong}]\geq\mathbb{E}[C_{\text{assist}_{i}}|\text{correct}]$ yields a $p$-value of $5.0\times 10^{-222}$, indicating that correct answers have greater consistency. We summarize this into Observation 3.
@@ 3.4 Model Consistency
In addition to fine-grained token consistency, it is crucial to quantify the trustworthiness of the model. Prior works typically assign model weights based on heuristics, such as uniform weighting Yao et al. (2024); Yu et al. (2024); Xu et al. (2024c) or prior knowledge Huang et al. (2024), or self-confidence metrics like perplexity Mavromatis et al. (2024); Liu et al. (2024). However, the inter-model consistency remains largely underexplored.
@@ Appendix C Datasets
Appendix D Limitations
@@ Appendix F Use Or Create Scientific Artifacts
Our work is built on public benchmarks and contributes new code resources to the community. For evaluation, we use widely adopted datasets including GSM8K, PIQA, SAMSum, TriviaQA, NQ, and MMLU, without making any changes to their original content. In addition, we introduce and release the codebase of CoRE (upon publication), providing a clear and well-structured repository to improve the accessibility of our research.
@@ F.2 Discuss The License For Artifacts
We comply with the licenses of all artifacts used or released in this work. The benchmarks are distributed under Creative Commons or public-domain terms, following the conditions specified by their original maintainers.
Model checkpoints retain their original open-source licenses, and our own code and generated data are released under the MIT license. All license files are included in our repository and described in the supplementary README.
@@ F.3 Artifact Use Consistent With Intended Use
F.4 Data Contains Personally Identifying Info Or Offensive Content
Although the benchmarks may contain public text with personal names or informal language, we ensure responsible use by:
  • Using only passages that are part of the officially released benchmark datasets.
  • Applying keyword-based filtering to remove or mask offensive content in analysis outputs.
F.5 Documentation Of Artifacts
We provide detailed documentation of all evaluation datasets and model combinations used in our experiments. Specifically, Appendix C of the appendix describes the coverage, task type, and evaluation metric for each dataset (e.g., reasoning, summarization, and knowledge QA), while Appendix H the LLMs included in the ensemble along with their sizes and sources. Together, these materials ensure transparency regarding the domains, data characteristics, and model diversity involved in our study.

### v2

@@ Abstract
Zhichen Zeng1thanks: Equal contribution., Qi Yu [^1] footnotemark: 1, Xiao Lin1, Ruizhong Qiu1, Xuying Ning1, Tianxin Wei1, Yuchen Yan1, Jingrui He1, Hanghang Tong1, 1University of Illinois Urbana-Champaign {zhichenz,qiyu6,htong}@illinois.edu
@@ 1 Introduction
Extensive efforts on test-time LLM ensemble can be broadly categorized into two categories: token-level and response-level ensemble. Token-level ensemble aligns and fuses the token probabilities of different LLMs at each decoding step Yu et al. (2024); Yao et al. (2024); Huang et al. (2024); Xu et al. (2024c), enabling fine-grained real-time correction for each token generation. Response-level ensemble offers a more coarse-level ensemble by selecting either a complete response Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024) or a span Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024a) from candidate outputs. Despite their success in improving ensemble quality, existing approaches largely overlook ensemble robustness against noisy or erroneous signals. For instance, incorrect token alignments can result in faulty probability fusion, while errors in model predictions may further compromise the correctness of ensemble outputs. Therefore, it is crucial for ensemble methods to detect and mitigate such potential errors during inference to ensure reliable performance.
@@
Figure 2: Ensemble performance across six benchmarks. We report the average performance of different ensemble methods with (CoRE) and without CoRE (Vanilla).
To bridge this gap, we propose a plug-and-play technique named CoRE to enhance both robustness and performance of LLM ensemble. Our preliminary observations reveal that ensemble failures are closely tied to large discrepancies between the token distributions of different LLMs. Motivated by this, we propose to harness both token and model consistency to achieve a robust ensemble. At the token level, significant disparities in output probabilities for a specific token indicate fine-grained uncertainty, often stemming from token space misalignment. To address this, we introduce token consistency, which measures the disparity between each model’s token probability and a reference probability, serving as a low-pass filter to amplify reliable tokens and suppress inconsistent ones. At the model level, large divergence among full token distributions reflects model conflicts. To capture this, we define model consistency that promotes model outputs with high self-confidence and low divergence from peers, thereby strengthening consistent models and downweighting unreliable ones. A key advantage of the proposed CoRE is that it is orthogonal to various token-level ensemble methods, and thus can be seamlessly integrated with no additional inference cost. We summarize the main contributions as follows.
  • We systematically analyze and improve the robustness of LLM ensembles.
  • We assess consistency at token and model levels to enhance ensemble performance and robustness. CoRE can be seamlessly integrated with various ensemble methods to enable test-time correction with no additional cost.
  • We conduct experiments across diverse benchmark tasks, model combinations, and ensemble methods. As shown in Figure 2, CoRE consistently improves baseline ensemble methods, achieving an average performance gain of 1.3% and 2.8% on Top-2 and Top-3 model ensemble, respectively.
@@ 2.1 Test-time LLM Ensemble
Response-level Ensemble selects the most promising response from the multiple outputs to ensemble LLMs at a coarser-grained level. One line of works select or synthesize a full response among model outputs by either training-free approaches, such as perplexity scoring or majority voting Jiang et al. (2023b); Lv et al. (2024a); Tekin et al. (2024), or training-based approaches that learn to rank responses Si et al. (2023). Another line of works Xu et al. (2024b); Lv et al. (2024b); Liu et al. (2024a); Cui et al. (2026) propose ensembling at the span level, aiming to strike a balance between fine-grained correction and context-aware decision-making.
@@ 2.2 Model Consistency
Multi-model Consistency combines outputs from different LLMs through majority voting Trad and Chehab (2024); Niimi (2025) or collaborative reasoning Wang et al. (2024); Liang et al. (2023); Li et al. (2024b). Despite the improved robustness, most focus on response-level agreement, overlooking fine-grained token-level consistency essential for detecting subtle errors in real time.
@@ 3 Methodology
In this section, we present our proposed CoRE for a trustworthy LLM ensemble. We begin by introducing the notation and formally defining the LLM ensemble problem in Section 3.1. Then, in Section 3.2, we reveal a strong correlation between ensemble performance and consistency scores at both the token and model levels. Motivated by these findings, we introduce our token-level consistency measure in Section 3.3 and our distribution-level consistency measure in Section 3.4, both designed to improve ensemble performance and reliability.
@@ 3.1 Problem Formulation
Formally, we are given a main model with vocabulary set $\mathcal{V}_{\text{main}}$ and $N$ assistant models with vocabulary set $\mathcal{V}_{\text{assist}_{i}}$ with $i\in\{1,2,...,N\}$. We denote the predicted token distributions of main and assistant models as $\bm{p}_{\text{main}}$ and $\bm{p}_{\text{assist}_{i}}$, respectively. LLM ensemble first learns a token alignment matrix $\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{assist}_{i}}|\times|\mathcal{V}_{\text{main}}|}$ between the main model and the $i$-th assistant model, where $\bm{A}_{i}(v,u)=1$ indicates that token $v\in\mathcal{V}_{\text{assist}_{i}}$ is aligned with token $u\in\mathcal{V}_{\text{main}}$. Token prediction probabilities can be further projected into the main model token space via $\bm{p}_{\text{assist}_{i}}\bm{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{\text{main}}|}$. For simplicity, we denote the aligned probability as $\tilde{\bm{p}}_{\text{assist}_{i}}=\bm{p}_{\text{assist}_{i}}\bm{A}_{i}$. The aligned probabilities can be further ensembled via
@@ 3.2 Observations
First, we examine how token disparity, measure by the difference between the aligned probability and the reference probability on a specific token $v$, i.e., $\bm{\delta}_{i}(v)=|\tilde{\bm{p}}_{\text{assist}_{i}}(v)-\bm{p}^{*}(v)|$, reflects the alignment correctness. We experiment on 100 randomly-sampled questions from NQ dataset, and as shown in Figure 4(a), aligned tokens concentrate at lower disparities, while misaligned tokens exhibit a heavier right tail. A one-sided test of $H_{0}:\mathbb{E}_{v\in\text{misalign}}\bm{\delta}_{i}(v)\leq\mathbb{E}_{v\in\text{align}}\bm{\delta}_{i}(v)$ yields a $p$-value of $2.7\times 10^{-44}$, indicating that aligned tokens indeed have smaller disparity. We summarize this in Observation 1.
@@
Secondly, we evaluate how model confidence, measured by the entropy of the token probability $H(\tilde{\bm{p}}_{\text{assist}_{i}})$, relates to the answer correctness. We experiment on PIQA dataset, and as shown in Figure 4(b), correct answers exhibit lower entropy than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[H(\tilde{\bm{p}})|\text{wrong}]\leq\mathbb{E}[H(\tilde{\bm{p}})|\text{correct}]$ yields a $p$-value of $7.7\times 10^{-108}$, confirming that wrong answers have larger entropy than correct ones. We summarize this into Observation 2.
@@
Thirdly, we evaluate how model consistency, measured by the sum of RBF-transformed token disparities, i.e., $C_{\text{assist}_{i}}(v)=\allowbreak\sum_{v\in\mathcal{V}_{\text{main}}}\exp(-\bm{\delta}_{\text{assist}_{i}}(v)/\sigma)$, varies between correct and wrong answers. We experiment on PIQA dataset, and as shown in Figure 4(c), correct answers exhibit higher model consistency than wrong ones. A one-sided test of $H_{0}:\mathbb{E}[C_{\text{assist}_{i}}|\text{wrong}]\geq\mathbb{E}[C_{\text{assist}_{i}}|\text{correct}]$ yields a $p$-value of $5.0\times 10^{-222}$, indicating that correct answers have greater consistency. We summarize this into Observation 3.
@@ 3.4 Model Consistency
In addition to fine-grained token consistency, it is crucial to quantify the trustworthiness of the model. Prior works typically assign model weights based on heuristics, such as uniform weighting Yao et al. (2024); Yu et al. (2024); Xu et al. (2024c) or prior knowledge Huang et al. (2024), or self-confidence metrics like perplexity Mavromatis et al. (2024); Liu et al. (2024a). However, the inter-model consistency remains largely underexplored.
@@ 4.4.1 Scaling Results
4.4.2 Ablation Study
On cosistency scores.
On entropy weight.
We ablate the entropy term in the model consistency in Eq. (3), and the results are presented in Figure 9. From these results, we draw two key observations. First, ablating the entropy term consistently leads to performance degradation across all baselines. As shown by the hatched regions in the figure, reintroducing the entropy term yields performance gains ranging from +0.27 to +1.28, effectively validating its contribution to the model’s overall accuracy. Notably, the UniTE model benefits most significantly, with a gain of +1.28. Second, even when the entropy term is ablated (colored bars), the performance remains robust compared to the vanilla ensemble, validating the effectiveness of the underlying consistency design independent of the entropy regularization.
Figure 9: Ablation study of the entropy term. Colored bars indicate model consistency without the entropy term, while hatched bars indicate model consistency with the entropy term (CoRE).
On token consistency design. The token consistency design in Eq. (2) is asymmetric, where only assist models are penalized while the main model remains original. To validate this design, we compare it against a symmetric variant where token consistency is applied to both main and assist models. The results are reported in Figure 10. We observe that the symmetric application (colored bars) consistently underperforms the proposed assistant-only design (hatched bars, CoRE). As indicated by the annotated gains in the figure, removing the penalty from the main model leads to significant improvements, such as +1.78 on PIQA with GAC and +1.66 on NQ with EVA. This confirms that the primary inconsistencies stem from assistant-to-main misalignment; therefore, penalizing the main model introduces unnecessary constraints that hurt the overall ensemble performance.
Figure 10: Ablation study of the token consistency. Colored bars indicate applying token consistency on both main and assist models, while hatched bars indicate only applying token consistency on assist models (CoRE).
4.5 Hyperparameter Study
Figure 11: Hyperparameter study on $\sigma$ in RBF function, $\alpha$ and $\beta$ in power function, and $\gamma$ in sigmoid function.
We conduct a sensitivity analysis on the hyperparameters governing the consistency functions: $\sigma$ in the RBF kernel, $\alpha$ and $\beta$ in the power function, and $\gamma$ in the sigmoid function. As shown in Figure 11, across a broad range of values, CoRE is fairly robust to these hyperparameters and does not require heavy fine-tuning.
Conceptually, these hyperparameters control the sharpness of the consistency score, i.e., how aggressively inconsistent tokens and models are penalized.High values of $\alpha,\beta,\gamma$ (or small $\sigma$) yield "peaked" scores that strictly downweight inconsistencies, effectively filtering noise when assistant models are weak. Conversely, lower values (or large $\sigma$) produce smoother scores that tolerate more diversity, which is beneficial when the ensemble consists of strong, reliable models.
@@ 5 Conclusion
Limitations
Acknowledgements
This work is supported by NSF (2433308) and AFOSR (FA9550-24-1-0002). The content of the information in this document does not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.
@@ Appendix C Datasets
Appendix D More Related Works
Pre-trained Foundation Models.
The landscape of NLP has been revolutionized by the emergence of LLMs trained on massive corpora Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023a), with wide applications in reasoning Lin et al. (2024a, 2025a), language understanding Team et al. (2023); Team (2024), and many more Li et al. (2025a, b). However, due to differences in training data, architectures, and tokenization schemes, these models exhibit diverse strengths and weaknesses across different tasks Wang et al. (2023); Team (2023). For instance, some models may excel at mathematical reasoning Guo et al. (2025); Team et al. (2023) while others specialize in dialogue or summarization Achiam et al. (2023). The heterogeneity among foundation models serves as the primary motivation for test-time ensemble strategies, which aim to integrate their complementary capabilities to achieve superior performance and reliability.
Model Ensemble and Alignment.
In the era of big data and AI, machine learning models are inherently heterogeneous Zheng et al. (2024); Zeng et al. (2025a, b), characterized by diverse model architectures Hochreiter and Schmidhuber (1997); Vaswani et al. (2017); Dosovitskiy (2020), data modalities Yan et al. (2021b, a, 2022); Zeng et al. (2023b, 2024b, 2025c), and distinct tokenization frameworks Sennrich et al. (2016); Kudo and Richardson (2018); Wu et al. (2016). Model ensemble is a well-established paradigm to combine various model capabilities for improved performance, with wide applications in multi-modal learning Zeng et al. (2026); Ning et al. (2025), language models Huang et al. (2024); Yao et al. (2024); Yu et al. (2024), time-series analysis Liu et al. (2025); Lin et al. (2024b, 2025b), recommendation Liu et al. (2024b); Liang et al. (2025); Yoo et al. (2024), social analysis Xu et al. (2024a); Yan et al. (2024b, a, 2025) and many more. A prerequisite for ensembling heterogeneous models is to bridge their gaps via alignment Zeng et al. (2023a, 2024a); Yu et al. (2025b, a). Specifically, token alignment ensures that probability distributions from heterogeneous LLMs are mapped onto a unified space before fusion. Early approaches align token spaces via minimum edit distance Fu et al. (2023); Wan et al. (2024); Mavromatis et al. (2024), which captures structural differences but often incurs high computational overhead. Subsequent methods focus on vocabulary mapping in the text space Yu et al. (2024); Yao et al. (2024) by constructing mappings based on exact or prefix matches between token strings. Recent works utilize token embeddings to exploit semantic information that text matching may fail to capture Huang et al. (2024); Xu et al. (2024c), learning projection functions or cross-model embeddings to transform heterogeneous token distributions into a shared representation space.
LLM Robustness.
Ensuring model robustness is critical for LLM deployment in real-world applications. Existing research largely focuses on safety against adversarial attacks and stability under distribution shifts. Adversarial robustness studies show that LLMs are vulnerable to jailbreak prompts Wei et al. (2023); Lin et al. (2026) or perturbations that trigger unintended behaviors Zou et al. (2023). Meanwhile, out-of-distribution (OOD) robustness aims to maintain performance amidst noisy inputs or domain shifts Zhu et al. (2023). Common mitigation techniques range from safety alignment (RLHF) Bai et al. (2022); Zhang et al. (2025) to test-time interventions like uncertainty quantification Kadavath et al. (2022); Lin et al. (2023). While most existing works focus on individual model, we study the multi-model setting, showing that consistency constraints can serve as a robustification mechanism to mitigate the impact of unreliable signals within an ensemble.
@@ Appendix F Use Or Create Scientific Artifacts
Our work is built on public benchmarks and contributes new code resources to the community. For evaluation, we use widely adopted datasets including GSM8K, PIQA, SAMSum, TriviaQA, NQ, and MMLU, without making any changes to their original content. In addition, we introduce and release the codebase of CoRE at https://github.com/zhichenz98/CoRE-EACL26, providing a clear and well-structured repository to improve the accessibility of our research.
@@ F.2 Discuss The License For Artifacts
We comply with the licenses of all artifacts used or released in this work. The benchmarks are distributed under Creative Commons or public-domain terms, following the conditions specified by their original maintainers. Model checkpoints retain their original open-source licenses, and our own code and generated data are released under the MIT license. We will release our code upon publication.
@@ F.3 Artifact Use Consistent With Intended Use
F.4 Documentation Of Artifacts
We provide detailed documentation of all evaluation datasets and model combinations used in our experiments. Specifically, Appendix C describes the coverage, task type, and evaluation metric for each dataset (e.g., reasoning, summarization, and knowledge QA), while Appendix H the LLMs included in the ensemble along with their sizes and sources. Together, these materials ensure transparency regarding the domains, data characteristics, and model diversity involved in our study.